A piece of rope is 108 cm long. One piece is three less tan twice the first. How long is each piece?
Do you set it up doing the 1st rope =, and 2nd rope = or just use Xs

Answer by checkley77(12844)

X+2X-3=108
3X=108+3
3X=111
X=111/3
X=37 CM. IS THE LENGTH OF THE FIRST PIECE OF ROPE.
2*37-3=74-3=71 CM. IS THE LENGTH OF THE SECOND PIECE OF ROPE.
PROOF:
37+71=108
108=108